Silver is a metal commonly used in jewelry and photography. It crystallizes with a FCC unit cell 0.407 nm on an edge. What is the theoretical density (g/mL) of silver in g/cm3 if its atomic weight is 107.87 g/mol?


1
Expert's answer
2021-10-04T22:30:01-0400

The number of atoms in the unit cell of an FCC structure , n=4

"a=0.407nm=0.407\times 10^{-7}cm"

Avogadro number, "N_A=6.022\times 10^{23}"

"M=atomic mass=107.87g\/mol."

density of silver "=\frac{n\times M}{N_A \times a^3}=\frac{4\times 107.87}{6.022\times 10^{23} \times (0.407\times 10^{-7})^3}=10.62g/cm^3"